OpenStudy (anonymous):

The width of a rectangle is 8 inches less than one-half of its length. If the perimeter of the rectangle is 38 inches, find the area of the rectangle.

OpenStudy (anonymous):

If the length is x, then the width would be (x/2-8) Perimeter = x + x +(x/2-8) + (x/2-8) 38=2x +x -16 Solve for x and you get x = 18 inches. Length = x = 18 inches Width = (x/2-8) = 1 inches Area of a rectangle is length times width Area =\[18\times1\]=18 inches squared.